# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ 2.4.2 A large deviation computation
In the following, we work with a sequence $\{{\bm{\xi}}_{j}=(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ of i.i.d. centered random variables in $\mathbb{R}^{d}$ that are not necessarily spherically symmetric but have densities belonging to $L^{r}(\mathbb{R}^{d})$ for some $r>1$ and a positive definite covariance matrix $V$. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel-Legendre transform. [^10] Not to confuse with the functions $\Lambda$ and $I$ for the first coordinate which is defined on $\mathbb{R}$, since we may assume $d\geqslant 2$. Let also $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})$. By definition,
@@
Fix a large constant $C>0$ and a dimension $d\geqslant 1$. Suppose that the $\mathbb{R}^{d}$-valued random variable ${\bm{\xi}}$ satisfies conditions (A2), (A5), and (A6) from Section 1.1. Let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as ${\bm{\xi}}$ and $(X,\mathbf{Y})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Fix $\overline{c}>\underline{c}>0$.
@@
    $\mathbb{P}(\mathbf{Y}\in B_{\mathbf{0}}(1)\mid X\geqslant x+c)\ll e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{0})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.$ (41)
@@
    $\begin{split}&\mathbb{P}(\mathbf{Y}\in B_{\mathbf{y}}(1)\mid X\in[x+c,x+c+a(n)])\\ &\hskip 56.9055pt\gg e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{y})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.\end{split}$ (42)
@@ 2.4.3 Upper bound of Theorem 3
$\widehat{P}_{n}:=\left\{v\in V_{n}:\leavevmode\nobreak\ \forall 0\leqslant k\leqslant n,\leavevmode\nobreak\ \bm{\eta}_{v,n}(k)-\frac{k\mathbf{x}}{n}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right\}$
@@
for some $C>0$ independent of $x$, which in turn gives the upper bound for $\tau_{x}$ by Theorem 2. We compute first and second moments for $\widehat{\Delta}_{n,x}$. Define
$\widehat{Q}_{n}:=\left\{v\in V_{n}:\leavevmode\nobreak\ \forall 0\leqslant k\leqslant n,\bm{\eta}_{v,n}(k)\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right\}.$
Applying the change of measure (71) and using a similar change of measure argument as in the proof of Proposition 17, we have
$\mathbb{E}[\widehat{\Delta}_{n,x}]$ $=\rho^{n}\mathbb{P}\left(v\in\widehat{P}_{n},\bm{\eta}_{v,n}(n)\in B_{x}\right)\gg\rho^{n}e^{-nI(x/n,\mathbf{0})}\mathbb{Q}\left(v\in\widehat{Q}_{n},\left\lVert\bm{\eta}_{v,n}(n)\right\rVert\leqslant 1\right).$
@@
$\mathbb{Q}\left(v\in\widehat{Q}_{n},\left\lVert\bm{\eta}_{v,n}(n)\right\rVert\leqslant 1\right)\asymp n^{-p_{\alpha,d}-d/2}\asymp x^{-p_{\alpha,d}-d/2}.$
@@
[…]
@@
×$\hskip 56.9055pt\times\left(e^{-sI({x}/{n},\mathbf{0})-\mathbf{c}_{2}\cdot(-\mathbf{y})}s^{-p_{\alpha,d}-d/2}(1+\left\lVert\mathbf{y}\right\rVert)^{2p_{\alpha,d}}\right)^{2}$
@@
where in the last step we argue similarly as in the proof of Lemma 2.7 of [13]. Combined with (47) and the Paley-Zygmund inequality leads to (46). This proves the desired upper bound.
@@ 2.4.4 Lower bound of Theorem 3
where $\bm{\psi}_{n}(k):=(C_{1}+C_{2}(\log(\min(k,n-k)))_{+})\,\mathbf{c}_{2}$ for some $C_{1},C_{2}>0$ to be determined, and for $10\leqslant\ell\leqslant n-$ 1,
@@
Next, we bound $\mathbb{E}[M_{x,\ell}]$ for $10\leqslant\ell\leqslant n-$ 1. We condition on the location of $\bm{\eta}_{v,n}(\ell)$, which may take value in ${\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$. This leads to
$\begin{split}\mathbb{E}[M_{x,\ell}]\leqslant\rho^{\ell}\sum_{\begin{subarray}{c}\mathbf{u}\in\mathbb{Z}^{d}\\ \mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\end{subarray}}&\mathbb{P}\left(\forall 0\leqslant k\leqslant\ell,\,\mathbf{S}_{k}\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),\,\mathbf{S}_{\ell}\in B_{\mathbf{u}}\right)\\ &\times\mathbb{P}\left({\bm{\xi}}\in-\mathbf{u}+\frac{(\ell+1)\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)+(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}\right).\end{split}$ (48)
@@
Inserting the above estimates back to (48) with the change of variable $\mathbf{v}=\mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))$ leads to
@@
×$\hskip 85.35826pt\times\exp\left(-(1+\delta)\inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\right)$
@@
×$\hskip 85.35826pt\times\sum_{j=0}^{\infty}e^{-(1+\delta)j}\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\\ \inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\in[j,j+1]\end{subarray}}\left\lVert\mathbf{v}\right\rVert e^{-\mathbf{c}_{2}\cdot\mathbf{v}},$
@@
The rest of the arguments is standard. Recall the weaker bound of (11). If a particle reaches $B_{x}$ at time $n-t$ where $t=O(\log x)$, the probability that one of its descendants is located in $B_{x}$ at time $n$ is $\gg(\log x)^{-d}$. On the other hand, $\mathbb{E}[M_{x}]\ll x^{-\varepsilon/10}=o((\log x)^{-d})$. As a consequence, with probability $o(1)$, there is no particle present in $B_{x}$ before time $n=\widehat{A}_{2}(x)$. This finishes the proof of the lower bound of Theorem 3.
@@ 2.5.1 Proof of Corollary 4
By Borel–Cantelli lemma,
[…]
@@ 2.5.2 Proof of Corollary 5
where $\tau_{x,1/n}$ denotes the first passage time to a ball of radius $1/n$ centered at $(x,\mathbf{0})$ and $C_{1}>0$. We need to show that $\widehat{c}_{1}$ is large in all directions. For a $d\times d$ orthogonal matrix $\Theta$, denote by $\widehat{c}_{1}(\Theta)$ the $\widehat{c}_{1}$ corresponding to the BRW with jump distribution $\Theta{\bm{\xi}}$. It follows from assumption (A5) that the domain of $I$ contains a neighborhood of $\mathbf{0}$, and hence
@@
for some $C_{3},C_{4}>0$. Thus $\sum_{n}\mathbb{P}(E_{n})<\infty$. The rest arguments follow in the same way as Corollary 2.5 of [43].
@@ 3.1 The delayed branching model
In this section, we revisit the applications of BRW to modeling polymer networks. In particular, the FPT for the BRW to reach a unit ball at a distance of $x$ can approximate the shortest path (SP) between cross-links placed at a distance $x$ away from each other in a polymer network. Experimentally, obtaining the SP statistics is infeasible. However, extracting the SP statistics from computational simulations of polymer networks using coarse-grained molecular dynamics (CGMD) has been shown promising [50, 51]. As argued in [54], the SP statistics can be reproduced by FPT statistics of a small variation of the classical BRW model (the reason of which will be explained below), which we refer to as the delayed branching model.
Consider the classical BRW model satisfying $p_{3}>0$ and $p_{0}+p_{1}+p_{3}=1$ (i.e., only branching instances are into three particles) with a spherically symmetric jump distribution (for simplicity). The choice of branching into three particles is to be consistent with the real polymer network being modeled where branching represents the cross-link (bond) between two polymer chains. The first particle would correspond to the continuation of the first polymer chain beyond the point of cross-link whereas, the other two particles would correspond to the segments of the second polymer chain (see [54] for a detailed description).
@@
Figure 4: BRW tree representations of polymer networks in the classical and the delayed branching regimes. The red path from $A$ to $A^{\prime\prime}$ represents the first polymer chain and the blue path from $B$ to $B^{\prime\prime}$ represents the second chain that cross-links with the first. In the BRW tree genealogy, the particles $A^{\prime\prime}$, $B$, and $B^{\prime\prime}$ are the descendants of the particle starting from $A$.
@@
Let $\widetilde{c}_{1}$ satisfy $I(\widetilde{c}_{1})=\log\tilde{\rho}$ where the rate function $I$ is defined in (1), and $\widetilde{c}_{2}=I^{\prime}(\widetilde{c}_{1})$. We define $\widetilde{A}(x)$ similarly as in (3) with $c_{1},\,c_{2}$ replaced by $\widetilde{c}_{1},\,\widetilde{c}_{2}$. The parameters $p_{3}$ and $p_{0}$ are denoted by $\widetilde{\kappa}$ and $\widetilde{\nu}$ in [54] respectively, where $\widetilde{\kappa}$ represents the branching rate of the BRW and $\widetilde{\nu}$ represents the death rate of the branches to account for the finite length of the polymer chains in the CGMD simulation. Let $\widetilde{\tau}_{x}$ be the first passage time to $B_{x}$.
@@
Theorem 18 partially resolves Conjecture 1 of [54]. We sketch its proof below, which will be very similar to the proof of Theorem 1. A key ingredient in the proof of Theorem 1 is the analysis of the expected number $N_{n}$ of particles present at time $n$, conditioned upon survival. If the branching events are independent across layers, then $N_{n}=\rho^{n}$. On the other hand, for the delayed branching BRW, due to the two-step branching regime, the branching events are not independent. The following two lemmas perform a similar analysis related to the number of particles for the delayed branching model. In particular, the analogue for our expected number of descendants for each particle, $\rho$, becomes $\tilde{\rho}$ for the delayed branching model.
Lemma 19 (Lemma 5 of [54]).
@@
Moreover, $\mathbb{E}[N_{u,s}{\mathbbm{1}}_{S}]$ can be controlled by the expected number of particles without conditioning upon survival, which is $\asymp\tilde{\rho}^{s}$. Otherwise, if $u$ initiates a branching sub-event of type II, an extra factor of $2$ exists but will be absorbed into the asymptotics. Therefore we conclude that $\mathbb{E}[N_{u,s}\mid S]\asymp\tilde{\rho}^{s}$ (that $\mathbb{P}(S)\in(0,1)$ has been proven in Lemma 5 of [54]), and hence $N_{n,s}\asymp\tilde{\rho}^{n-s}(\tilde{\rho}^{s})^{2}=\tilde{\rho}^{n+s}$. ∎
We briefly identify the modifications required for the proof of Theorem 18. Note that the only difference between the classical BRW and delayed branching BRW is the branching structure, i.e., the underlying tree that describes the genealogy of the particles. Given that Theorem 2 holds, the proof of Theorem 1 depends on such structure only through (17) and (18). The corresponding modifications can be made by replacing $\rho$ by $\tilde{\rho}$ in view of Lemmas 19 and 20. The proof of Theorem 2 depends on large deviation estimates for the maximum of delayed branching BRW in (28). Nevertheless, this applies by performing the same proof as in [29] by comparing against the maximum of independent random walks (Theorems 3.1 and 5.2 therein). The other changes needed in Theorem 2 are minor and will be omitted for brevity, as the constants can always be adjusted to take into account the delayed branching property. Otherwise, the proof of Theorem 18 is almost verbatim.

### v2

@@ 2.4.2 A large deviation computation
In the following, we work with a sequence $\{{\bm{\xi}}_{j}=(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ of i.i.d. centered random variables in $\mathbb{R}^{d}$ that are not necessarily spherically symmetric but have densities belonging to $L^{r}(\mathbb{R}^{d})$ for some $r>1$ and a positive definite covariance matrix $V$. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel–Legendre transform. [^11] Not to be confused with the functions $\Lambda$ and II for the first coordinate which is defined on $\mathbb{R}$, since we may assume $d\geqslant 2$. Let also $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})$. By definition,
@@
Fix a large constant $C>0$ and a dimension $d\geqslant 1$. Suppose that the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$} satisfies conditions (A2), (A5), and (A6) from Section 1.1. Let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$} and $(F_{n},\mathbf{J}_{n})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Fix $\overline{c}>\underline{c}>0$.
@@
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\mid F_{n}\geqslant x+c)\ll e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{0})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.$ (41)
@@
    $\begin{split}&\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1)\mid F_{n}\in[x+c,x+c+a(n)])\\ &\hskip 56.9055pt\gg e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{y})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.\end{split}$ (42)
@@ 2.4.3 Upper bound of Theorem 3
$\widehat{P}_{n}:=\left\{v\in V_{n}:\penalty 10000\ \forall 0\leqslant k\leqslant n,\penalty 10000\ \bm{\eta}_{v,n}(k)-\frac{k\mathbf{x}}{n}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right\}$
@@
for some $C>0$ independent of $x$, which in turn gives the upper bound for $\tau_{x}$ by Theorem 2. We compute first and second moments for $\widehat{\Delta}_{n,x}$. Applying the many-to-one formula, the change of measure (72), and using a similar change of measure argument as in the proof of Proposition 17, we have
$\mathbb{E}[\widehat{\Delta}_{n,x}]$ $=\rho^{n}\mathbb{P}\left(\forall 0\leqslant k\leqslant n,\penalty 10000\ \mathbf{S}_{k}-\frac{k\mathbf{x}}{n}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\mathbf{S}_{n}\in B_{x}\right)$
$\gg\rho^{n}e^{-nI(x/n,\mathbf{0})}\mathbb{Q}\left(\forall 0\leqslant k\leqslant n,\mathbf{S}_{k}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\left\lVert\mathbf{S}_{n}\right\rVert\leqslant 1\right).$
@@
$\mathbb{Q}\left(\forall 0\leqslant k\leqslant n,\mathbf{S}_{k}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\left\lVert\mathbf{S}_{n}\right\rVert\leqslant 1\right)\asymp n^{-p_{\alpha,d}-d/2}\asymp x^{-p_{\alpha,d}-d/2}.$
@@
Next, we compute the second moment of $\widehat{\Delta}_{n,x}$, following the same type of arguments that lead to (19). Recall the definition of $\widehat{H}_{v,n}(x)$ from (45). By conditioning on the closest integer point to $\bm{\eta}_{v,n}(n-s)\in(n-s)\mathbf{x}/n+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$ and applying (36) of Lemma 14, $\text{ for }v\sim_{s}w$ (recall that this means that the distance of $v$ and $w$ in the genealogical tree is equal to $2s$),
@@
$\hskip 56.9055pt\times\left(e^{-sI({x}/{n},\mathbf{0})-\mathbf{c}_{2}\cdot(-\mathbf{y})}s^{-p_{\alpha,d}-d/2}(1+\left\lVert\mathbf{y}\right\rVert)^{2p_{\alpha,d}}\right)^{2}$
@@
where in the last step we argue similarly as in the proof of Lemma 2.7 of [bramson2016convergence]. Combined with (47) and the Paley-Zygmund inequality leads to (46). This proves the desired upper bound.
@@ 2.4.4 Lower bound of Theorem 3
where $\bm{\psi}_{n}(k):=(C_{1}+C_{2}(\log(\min(k,n-k)))_{+})\,\mathbf{c}_{2}$ for some $C_{1},C_{2}>0$ to be determined, and for $0\leqslant\ell\leqslant n-1$,
@@
Next, we bound $\mathbb{E}[M_{x,\ell}]$ for $0\leqslant\ell\leqslant n-1$. We condition on the location of $\bm{\eta}_{v,n}(\ell)$, which may take value in ${\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$. This leads to
$\begin{split}&\hskip 14.22636pt\mathbb{E}[M_{x,\ell}]\\ &\leqslant\rho^{\ell}\sum_{\begin{subarray}{c}\mathbf{u}\in\mathbb{Z}^{d}\\ \mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\end{subarray}}\mathbb{P}\left(\forall 0\leqslant k\leqslant\ell,\,\mathbf{S}_{k}\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),\,\mathbf{S}_{\ell}\in B_{\mathbf{u}}\right)\\ &\hskip 128.0374pt\times\mathbb{P}\left({\bm{\xi}}\in-\mathbf{u}+\frac{(\ell+1)\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)+(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}\right).\end{split}$ (48)
@@
Inserting the above estimates back to (LABEL:eq:Mxl) with the change of variable $\mathbf{v}=\mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))$ leads to
@@
$\hskip 85.35826pt\times\exp\left(-(1+\delta)\inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\right)$
@@
$\hskip 85.35826pt\times\sum_{j=0}^{\infty}e^{-(1+\delta)j}\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\\ \inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\in[j,j+1]\end{subarray}}\left\lVert\mathbf{v}\right\rVert e^{-\mathbf{c}_{2}\cdot\mathbf{v}},$
@@
The rest of the arguments is standard. Recall the weaker bound of (11). If a particle reaches $B_{x}$ at time ntn-t where $t=O(\log x)$, the probability that one of its descendants is located in $B_{x}$ at time $n$ is $\gg(\log x)^{-d}$. On the other hand, $\mathbb{E}[M_{x}]\ll x^{-\varepsilon/10}=o((\log x)^{-d})$. As a consequence, with probability $o(1)$, there is no particle present in $B_{x}$ before time $n=\widehat{A}_{2}(x)$. This finishes the proof of the lower bound of Theorem 3.
@@ 2.5.1 Proof of Corollary 4
By the Borel–Cantelli lemma,
$\frac{\tau_{n}(2)}{n}\to\frac{1}{\widehat{c}_{1}}\quad\text{ and }\quad\frac{\tau_{n/2}(1/2)}{n/2}\to\frac{1}{\widehat{c}_{1}}\quad\text{a.s.}$
@@ 2.5.2 Proof of Corollary 5
where $\tau_{x,1/n}$ denotes the first passage time to a ball of radius $1/n$ centered at $(x,\mathbf{0})$ and $C_{1}>0$. We need to show that $\widehat{c}_{1}$ is large in all directions. For a $d\times d$ orthogonal matrix $\Theta$, denote by $\widehat{c}_{1}(\Theta)$ the $\widehat{c}_{1}$ corresponding to the BRW with jump distribution $\Theta{\bm{\xi}}$. By assumption (A5) and our assumption that $\phi_{\bm{\xi}}$ is well-defined in a neighborhood of $\mathbf{0}$, the domain of $I$ contains a neighborhood of $\mathbf{0}$, and hence
@@
for some $C_{3},C_{4}>0$. Thus $\sum_{n}\mathbb{P}(E_{n})<\infty$. The remaining arguments follow in the same way as Corollary 2.5 of [oz2023].
@@ 3.1 The delayed branching model
In this section, we revisit the applications of BRW in modeling polymer networks. A polymer network is a three-dimensional disordered structure consisting of long sequences of monomers, called polymer chains, connected via chemical bonds. It turns out that in the study of polymers, random walks provide a decent statistical approximation to polymer chains [flory1960elasticity]. In turn, any polymer chain can be bonded to another chain at a certain site (e.g. a monomer) also by chemical bonds called cross-links.
The resulting network topology is generally random. A chain segment is a connected segment of monomers inside a polymer chain with the property that each of its two ends is connected, via cross-links, to another polymer chain (not necessarily the same one), as mentioned earlier. Chain segments have varying (random) lengths.
As it turns out, macroscopic mechanical properties, such as elasticity and fracture toughness, are largely determined by the distribution of shortest paths (SP) between distant monomers [yin2020topological]. Understanding the statistics of these SP distributions is, therefore, central to predicting network behavior under deformation.
The FPT for the BRW to reach a unit ball at a distance of $x$ can approximate the SP between monomers placed at a distance $x$ from each other in a polymer network [zhang2024modeling]. The intuition can be appreciated from the picture in Figure 4. Locally, the polymer network at a given cross-link is approximated by the branching mechanism of the BRW.
In physical experiments, obtaining SP statistics is infeasible. However, extracting SP statistics from computational simulations of polymer networks using coarse-grained molecular dynamics (CGMD) has been shown promising [mohanty2025strength, yin2020topological, yin2024network]. Now, even though the BRW intuition is local, as argued earlier, and the SP distribution is a global object (involving long distances relative to the size of a single monomer), it was shown based on comparisons to CGMD simulations in [zhang2024modeling], that the SP statistics can be reproduced by FPT statistics of a small variation of the classical BRW model (the reason of which will be explained below), which we refer to as the delayed branching model. The reason for this small variation is to model explicitly the cross-link representing the segment that joins any two polymer chains.
Consider the classical BRW model satisfying $p_{3}>0$ and $p_{0}+p_{1}+p_{3}=1$ (i.e., only branching instances are into three particles) with a spherically symmetric jump distribution (for simplicity). The choice of branching into three particles is to be consistent with the real polymer network being modeled, where branching represents the cross-link (bond) between two polymer chains. The first particle would correspond to the continuation of the first polymer chain beyond the point of cross-link whereas, the other two particles would correspond to the segments of the second polymer chain (see [zhang2024modeling] for a detailed description).
@@
Figure 4: BRW tree representations of polymer networks in the classical and the delayed branching regimes. The red path from AA to $A^{\prime\prime}$ represents the first polymer chain and the blue path from $B$ to $B^{\prime\prime}$ represents the second chain that cross-links with the first. In the BRW tree genealogy, the particles $A^{\prime\prime}$, $B$, and $B^{\prime\prime}$ are the descendants of the particle starting from $A$.
@@
Let $\widetilde{c}_{1}$ satisfy $I(\widetilde{c}_{1})=\log\tilde{\rho}$ where the rate function $I$ is defined in (1), and $\widetilde{c}_{2}=I^{\prime}(\widetilde{c}_{1})$. We define $\widetilde{A}(x)$ similarly as in (3) with $c_{1},\,c_{2}$ replaced by $\widetilde{c}_{1},\,\widetilde{c}_{2}$. The parameters $p_{3}$ and $p_{0}$ are denoted by $\widetilde{\kappa}$ and $\widetilde{\nu}$ in [zhang2024modeling] respectively, where $\widetilde{\kappa}$ represents the branching rate of the BRW and $\widetilde{\nu}$ represents the death rate of the branches to account for the finite length of the polymer chains in the CGMD simulation. Let $\widetilde{\tau}_{x}$ be the first passage time to $B_{x}$.
@@
Theorem 18 partially resolves Conjecture 1 of [zhang2024modeling]. We sketch its proof below, which will be very similar to the proof of Theorem 1. A key ingredient in the proof of Theorem 1 is the analysis of the expected number $N_{n}$ of particles present at time $n$, conditioned upon survival. If the branching events are independent across layers, then $N_{n}=\rho^{n}$. On the other hand, for the delayed branching BRW, due to the two-step branching regime, the branching events are not independent. The following two lemmas perform a similar analysis related to the number of particles for the delayed branching model. In particular, the analogue for our expected number of descendants for each particle, $\rho$, becomes $\tilde{\rho}$ for the delayed branching model.
Lemma 19 (Lemma 5 of [zhang2024modeling]).
@@
Moreover, $\mathbb{E}[N_{u,s}{\mathbbm{1}}_{S}]$ can be controlled by the expected number of particles without conditioning upon survival, which is $\asymp\tilde{\rho}^{s}$. Otherwise, if $u$ initiates a branching sub-event of type II, an extra factor of $2$ exists but will be absorbed into the asymptotics. Therefore we conclude that $\mathbb{E}[N_{u,s}\mid S]\asymp\tilde{\rho}^{s}$ (that $\mathbb{P}(S)\in(0,1)$ has been proven in Lemma 5 of [zhang2024modeling]), and hence $N_{n,s}\asymp\tilde{\rho}^{n-s}(\tilde{\rho}^{s})^{2}=\tilde{\rho}^{n+s}$. ∎
We briefly identify the modifications required for the proof of Theorem 18. Note that the only difference between the classical BRW and delayed branching BRW is the branching structure, i.e., the underlying tree that describes the genealogy of the particles. Given that Theorem 2 holds, the proof of Theorem 1 depends on such structure only through (17) and (18). The corresponding modifications can be made by replacing $\rho$ by $\tilde{\rho}$ in view of Lemmas 19 and 20. The proof of Theorem 2 depends on large deviation estimates for the maximum of delayed branching BRW in (28). Nevertheless, this applies by performing the same proof as in [gantert2018large] by comparing against the maximum of independent random walks (Theorems 3.1 and 5.2 therein). The other changes needed in Theorem 2 are minor and will be omitted for brevity, as the constants can always be adjusted to take into account the delayed branching property. Otherwise, the proof of Theorem 18 goes through almost verbatim.